# HyperLoader: Integrating Hypernetwork-Based LoRA and Adapter Layers into Multi-Task Transformers for Sequence Labelling

### v2

@@ 1 Introduction
Although these methods considerably reduce the number of parameters to achieve good performance on different tasks, a specialized model for each task remains necessary. Multi-task learning reduces the computational cost by training a single model while enabling information sharing through capturing the common structure underlying all tasks. Recently, researchers have developed approaches combining parameter-efficient fine-tuning techniques in multi-task settings Stickland and Murray (2019); Pfeiffer et al. (2020, 2021); Rücklé et al. (2021). Nevertheless, this approach can lead to underfitting in high-resource tasks and overfitting in low-resource tasks Lee et al. (2017). Another potential issue is task interference, where an improvement in the performance of one task reduces the performance of other tasks Wang et al. (2019).
@@
Our model achieves the best average performance using the complete training and validation data for each task, as well as in a low-resource configuration with only 10% and 20% of the training and validation data available. We empirically demonstrate that the improvement is not only due to adding more trainable parameters but also by combining different parameter-efficient methods in a multi-task setting through hypernetworks, which adequately supports Sequence Labelling tasks in high- and low-resource scenarios.
@@ 2 Related Work
Parameter-efficient fine-tuning techniques are an alternative to full fine-tuning by updating a small number of parameters per task, yielding a performance similar to that of full fine-tuning. One of the most popular approaches in this area is the adapter Houlsby et al. (2019), a small trainable bottleneck layer added in each transformer block of a model. It consists of down and up projections and has shown similar performance to full-fine-tuning. Other popular parameter-efficient fine-tuning approaches are Prefix-tuning Li and Liang (2021) and LoRA Hu et al. (2021). Prefix-tuning prepends a fixed-length, learnable sequence of prefix tokens to the input embeddings while the original model parameters remain unchanged. This approach optimizes a continuous prompt, effectively guiding the model’s behaviour for specific tasks with minimal computational overhead. LoRA adds trainable low-rank matrices and combines their outputs with the original matrices in the self-attention layer of the transformer; the main difference with the previous approaches is that it does not employ any activation function.
@@
The HyperFormer++ model Karimi Mahabadi et al. (2021b) is the main foundation of our model. It uses task-conditioned hypernetworks to generate the weights of adapter layers and layer normalizations in a multi-task setting. The hypernetwork is trained to create adapter parameters specific to each task and layer, based on the embeddings of task and layer IDs. This hypernetwork is trained simultaneously for all tasks, allowing it to share knowledge across different tasks. At the same time, it reduces negative interference by producing distinct adapter layers for each individual task. The T5 model is used as the underlying model for all the experiments.
@@ 3 Methodology
In this section, we describe our approach in detail. First, we provide preliminary information by defining multi-task learning, which is the foundation of our work. Second, we describe the main components of our proposed HyperLoader model. Finally, we describe the format to convert Sequence Labelling into a sequence-to-sequence task.
@@ 3.6 SentT’ format
Converting Sequence Labelling into a sequence-to-sequence task is essential to using an encoder-decoder model. We use the SentT’ Farina et al. (2023) format for the input and output sequences to achieve this goal. This format is obtained by turning a labelled sequence $S$ into a sequence of tokens $x_{1}\dots x_{L}$ of size $L$ using white-space splitting and then interleaving these tokens with the spacial sentinel tokens used to pre-train T5 to obtain the input string $S_{in}=s_{0}x_{1}s1\dots x_{L}s_{L}$. We use the simplified Beginning, Inside and Outside (sBIO) format to represent the output strings, where given a set $T$ of labels to annotate a text span, $t\in T$ is used to represent any token that starts a labelled span, a single tag $I$ for each token that continues a labelled span and $O$ to tag tokens that do not belong to labelled spans. Analogous to the input sequence, we interleave the tokenized output string with the sentinel tokens to obtain the output sequence using the SentT’ format $S_{out}=s_{0}t_{1}s_{1}\dots t_{L}s_{L}$. Table 1 shows an example from the SNIPS dataset of how data is converted into the SentT’ format.
Original text
play the song little robin redbreast
@@
Table 1: Example of an input/output instance transformed into the SentT’ format. The text is tokenized using white-space splitting and then interleaving these tokens with the spacial sentinel tokens used to pre-train T5 to obtain the input string. Notice the use of the sBIO format to represent the output, where a label $t\in T$ is used to represent any token that starts a labelled span, a single tag $I$ for each token that continues a labelled span and $O$ to tag tokens that do not belong to labelled spans.
@@ 4 Experiments
This section describes our experiments for Sequence Labelling using the HyperLoader model. First, we introduce the datasets we used to train and evaluate our approach. Second, we provide the implementation, hyper-parameter and evaluation details of all experiments.
@@ 4.1 Corpora
We use seven publicly available corpora for Sequence Labelling tasks to cover different domains and distributions and gather robust empirical evidence. For slot-filling, we use four different dialogue-oriented datasets: ATIS Hemphill et al. (1990), SNIPS Coucke et al. (2018), and the English portions of two multilingual datasets, mTOP Li et al. (2021) and mTOP Schuster et al. (2019). For Named Entity Recognition, we use the MIT Corpora (Movie, MovieTrivia and Restaurant) [^1] Downloaded from https://groups.csail.mit.edu/sls/.
@@
Table 2: Statistics per partition of the used datasets. We use the training and validation partitions of Goo et al. (2018) for the ATIS and SNIPS dataset. For the MIT Corpora, we manually create the training and validation partitions following previous work Raman et al. (2022); Farina et al. (2023). Finally, we use the original training and validation partitions for the English portions of mTOP and mTOD.
4.2 Experimental details
We perform experiments to compare our model with other parameter-efficient fine-tuning approaches. The first goal is to observe the benefits of combining different parameter-efficient methods in a multi-task setting compared to a single-task setting. The second goal is to observe the benefits of the proposed model compared to other hypernetwork-based approaches in a multi-task setting.
We compare our approaches with the UniPELT framework Mao et al. (2022), HyperFormer model Karimi Mahabadi et al. (2021b) and Hyperdecoder model Ivison and Peters (2022) described in section 2. For all the implementations, we use the HuggingFace’s Transformers Wolf et al. (2020) and AdapterHub Poth et al. (2023) libraries.
@@ 5.1 Full dataset performance
Table 3 shows the performance of our proposed approach and the baselines using the 100% of training and evaluation data. The UniPELT framework follows a single-task fine-tuning approach, therefore, the trainable parameters correspond to each dataset considered. For this framework, we report the results using BERT as in the original paper and T5 to compare the performance depending on the model’s architecture. Initial experiments show that the use of the Prefix-tuning method and the gating mechanism decreases the performance of the encoder-decoder model; for this reason, we do not use these elements when applying UniPELT to T5.
@@
Based on the results in Table 3, it is possible to observe that our proposed HyperLoader model outperforms the baselines in three of the seven datasets and obtains the highest average with 0.8811, followed by the HyperFormer model (0.8795), UniPELT with T5 (0.8740), the Hyperdecoder model (0.8709) and UniPELT with BERT (0.8537). In terms of single-task and multi-task settings, it is possible to observe that the combination of UNiPELT with T5 outperforms the other models in only one dataset (ATIS, with 0.9611); this result demonstrates that the use of hypernetworks improves the sharing information capacity of the models while reducing the negative transfer between tasks. Finally, when comparing the multi-task approaches based on hypernetworks, HyperLoader obtained a better average result than the HyperFormer model, which is the primary foundation of our work. In addition, since Hyperdecoder has more trainable parameters, the results indicate that the superior performance of our model is not only due to having more trainable parameters than the HyperFormer approach but also due to the combination of different parameter-efficient methods in a multi-task setting.
UniPELT UniPELT HyperFormer HyperLoader Hyperdecoder
Dataset BERT[110M/1.8M] $\clubsuit$ T5[220M/4.5M] $\clubsuit$ T5[228M/5M] T5[229M/6M] T5[239M/16M]
MovieTriva 0.6873 0.7210 0.7295 0.7140 0.7072
Movie 0.8500 0.8839 0.8788 0.8745 0.8700
Restaurant 0.7689 0.8039 0.8136 0.7965 0.7891
ATIS 0.9446 0.9611 0.9595 0.9599 0.9519
SNIPS 0.9238 0.9478 0.9462 0.9592 0.9377
mTOP 0.8432 0.8374 0.8638 0.8983 0.8752
mTOD 0.9580 0.9627 0.9648 0.9652 0.9649
Average 0.8537 0.8740 0.8795 0.8811 0.8709
@@
Table 3: Performance of the HyperLoader model and the baselines using all train and evaluation examples. For each method, we indicate the used model and the total number of parameters / trainable parameters in brackets. $\clubsuit$ indicates a single-task fine-tuning approach; therefore, the number of trainable parameters is for each considered dataset. Bold fonts indicate the best result in each dataset and on average. Micro-averaged F1-score is reported.
@@ 5.2 Low-resource Setting Results
We evaluate the performance of our proposed model and the baselines in a low-resource setting to measure its robustness with a data scarcity constraint, a common scenario in NLP Hedderich et al. (2021). For this purpose, we down-sampled the training and validation partitions of all datasets to 10% and 20% of their original size and evaluated them in the full test subset. Table 4 shows the results of these experiments.
@@
The first part of Table 4 contains the performance of the HyperLoader model and the baselines when only 10% of the training and validation data is used. It is evident that the combination of the UniPELT framework and the BERT model experiences the most significant performance decrease, from an average of 0.8537 to 0.6795. In contrast, the T5-based approaches show a smaller performance decrease in both single-task and multi-task settings. These results suggest that an encoder-decoder model is more capable of handling low-resource settings in both single-task and multi-task configurations. For the multi-task approaches, our HyperLoader model outperforms the HyperFormer and Hyperdecoder approaches in six out of seven datasets and on the average performance with a score of 0.8331, followed by HyperFormer (0.8262) and Hyperdecoder (0.8161).
@@
The second part of Table 4 shows the results when 20% of the training and evaluation data is used. Using this amount of data our HyperLoader model outperforms the baselines in four of seven datasets and on the average performance. It is also noticeable that the combination of UniPELT and T5 using a single-task fine-tuning approach outperforms the other models only in the MovieTrivia dataset, with a micro-averaged F1-score of 0.7036 followed by HyperLoader with 0.6863, Hyperdecoder with 0.6767, HyperFormer with 0.6729 and finally UniPELT using BERT with 0.6105.
These results suggest that combining different parameter-efficient methods in a multi-task setting and generating their weights based on the task, the layer of the transformer model, and the position where the parameter-efficient method is applied, that is, the position of the Adapter layer or the LoRA matrix, is an effective method for Sequence Labelling tasks, yielding good performance using a parameter-efficient approach.
UniPELT UniPELT HyperFormer HyperLoader Hyperdecoder
Dataset BERT[110M/1.8M] $\clubsuit$ T5[220M/4.5M] $\clubsuit$ T5[228M/5M] T5[229M/6M] T5[239M/16M]
10% of Training and validation data
MovieTriva 0.5152 0.6717 0.6589 0.6695 0.6597
Movie 0.7809 0.8244 0.8311 0.8368 0.8202
Restaurant 0.5907 0.7049 0.7389 0.7499 0.7216
ATIS 0.5901 0.8919 0.9162 0.9172 0.9193
SNIPS 0.7398 0.8991 0.8943 0.8970 0.8727
mTOP 0.6048 0.7467 0.7976 0.8114 0.7760
mTOD 0.9351 0.9436 0.9467 0.9500 0.9436
Average 0.6795 0.8117 0.8262 0.8331 0.8162
20% of Training and validation data
MovieTriva 0.6105 0.7036 0.6729 0.6863 0.6767
Movie 0.8047 0.8444 0.8407 0.8546 0.8377
Restaurant 0.7114 0.7562 0.7517 0.7601 0.7401
ATIS 0.8048 0.9281 0.9382 0.9401 0.9311
SNIPS 0.8474 0.9235 0.9252 0.9238 0.9044
mTOP 0.7052 0.7863 0.8399 0.8474 0.8140
mTOD 0.9433 0.9522 0.9543 0.9541 0.9529
Average 0.7753 0.8420 0.8461 0.8523 0.8367
Table 4: Performance of the HyperLoader model and the baselines using 10% and 20% of the train and evaluation data. For each method, we indicate the used model and the total number of parameters / trainable parameters in brackets. $\clubsuit$ indicates a single-task fine-tuning approach; therefore, the number of trainable parameters is for each considered dataset. Bold fonts indicate the best result in each dataset and on average. Micro-averaged F1-score is reported.
Figure 2 illustrates the relationship between average performance and the percentage of trainable parameters per model. The symbol $\clubsuit$ denotes a single fine-tuning approach. It is important to note that, despite our model requiring more trainable parameters (6M) than the combination of UniPELT with T5 (4.5M), the percentage of parameters remains constant with increased training datasets.
Our proposed approach consistently outperforms other methods, using 10%, 20%, and 100% of the training and validation data. These results provide empirical evidence that combining different parameter-efficient fine-tuning methods in a multi-task setting is a robust strategy for low-resource scenarios. The performance improvement cannot be solely attributed to adding more parameters, as demonstrated by the hyperdecoder model having a more significant percentage of trainable parameters. Thus, the combination of parameter-efficient methods and weight generation emerges as a potent alternative for addressing Sequence Labelling tasks.
@@
Figure 2: Average performance-percentage of trainable parameters plot using different portions of the datasets.$\clubsuit$ indicates a single-task fine-tuning approach

### v3

@@ 1 Introduction
Although these methods considerably reduce the number of parameters to achieve good performance on different tasks, a specialized model for each task remains necessary. Multi-task learning reduces the computational cost by training a single model while enabling information sharing by capturing the common structure underlying all tasks. Recently, researchers have developed approaches combining parameter-efficient fine-tuning techniques in multi-task settings Stickland and Murray (2019); Pfeiffer et al. (2020, 2021); Rücklé et al. (2021). Nevertheless, this approach can lead to underfitting in high-resource tasks and overfitting in low-resource tasks Lee et al. (2017). Another potential issue is task interference, where an improvement in the performance of one task reduces the performance of other tasks Wang et al. (2019).
@@
Our model achieves the best average performance using the complete training and validation data for each task, as well as in a low-resource configuration with only 10% and 20% of the training and validation data available. We empirically demonstrate that the improvement is not only due to adding more trainable parameters but also by combining different parameter-efficient methods in a multi-task setting using hypernetworks, which adequately supports Sequence Labelling tasks in high- and low-resource scenarios.
@@ 2 Related Work
Parameter-efficient fine-tuning techniques are an alternative to full fine-tuning by updating a small number of parameters per task, yielding similar performance to full fine-tuning. One of the most popular approaches in this area is the adapter Houlsby et al. (2019), a small trainable bottleneck layer added in each transformer block of a model. It consists of down and up projections and has shown similar performance to full-fine-tuning. Other popular parameter-efficient fine-tuning approaches are Prefix-tuning Li and Liang (2021) and LoRA Hu et al. (2021). Prefix-tuning prepends a fixed-length, learnable sequence of prefix tokens to the input embeddings while the original model parameters remain unchanged. This approach optimizes a continuous prompt, effectively guiding the model’s behaviour for specific tasks with minimal computational overhead. LoRA adds trainable low-rank matrices and combines their outputs with the original matrices in the self-attention layer of the transformer; the main difference with the previous approaches is that it does not employ any activation function.
@@
The HyperFormer++ model Karimi Mahabadi et al. (2021b) is the main foundation of our model. It uses task-conditioned hypernetworks to generate the weights of adapter layers and layer normalizations in a multi-task setting. The hypernetwork is trained to create adapter parameters specific to each task and layer, based on the embeddings of task and layer IDs. This hypernetwork is trained simultaneously for all tasks, allowing it to share knowledge across different tasks. At the same time, it reduces negative interference by producing distinct adapter layers for each task. The T5 model is used as the backbone model for all the experiments.
@@ 3 Methodology
In this section, we describe our approach. First, we provide preliminary information by defining multi-task learning, which is the foundation of our work. Second, we describe the main components of our proposed HyperLoader model. Finally, we describe the format for converting Sequence Labelling into a sequence-to-sequence task.
@@ 3.5 HyperLoader: Task Conditional Hypernetworks
We use five hypernetworks for the HyperLoader model. For the Adapters, we consider two hypernetworks: one for the down projection and one for the up projection. Since we are using a reduction factor $r=32$ and the hidden dimension of the model is $d=768$, the down sample size is $\lfloor 768/32\rfloor=24$, the dimensions of the output matrices of the down and up projections are $768\times 24$ and $24\times 768$, respectively. We use two hypernetworks to generate the weights of the LoRA Matrices. One for the low-rank matrix $A$ and one for the low-rank matrix $B$. We use a low-dimensional rank $r=8$ for the Query and Value matrices in the self-attention and the cross-attention; therefore, the output matrices A and B dimensions are $768\times 8$ and $8\times 768$, respectively. Finally, we use one hypernetwork for the task-conditioned layer normalization that generates vector weights of $768$ dimensions.
@@ 3.6 SentT’ Format
Converting Sequence Labelling into a sequence-to-sequence task is essential to using an encoder-decoder model. We use the SentT’ Farina et al. (2023) format for the input and output sequences to achieve this goal. This format is obtained by turning a labelled sequence $S$ into a sequence of tokens $x_{1}\dots x_{L}$ of size $L$ using white-space splitting and then interleaving these tokens with the spacial sentinel tokens used to pre-train T5 to obtain the input string $S_{in}=s_{0}x_{1}s1\dots x_{L}s_{L}$. We use the simplified Beginning, Inside and Outside (sBIO) format to represent the output strings, where given a set $T$ of labels to annotate a text span, $t\in T$ is used to represent any token that starts a labelled span, a single tag $I$ for each token that continues a labelled span and $O$ to tag tokens that do not belong to labelled spans. Analogous to the input sequence, we interleave the tokenized output string with the sentinel tokens to obtain the output sequence using the SentT’ format $S_{out}=s_{0}t_{1}s_{1}\dots t_{L}s_{L}$. Table 1 illustrates how to convert the text "play the song little robin redbreast" from the SNIPS dataset into the SentT’ format.
@@
Table 1: Example of an input/output instance transformed into the SentT’ format. The text is tokenized with white-space splitting and then interleaved with sentinel tokens used for T5 pre-training to obtain the input string. The output uses the sBIO format: a label $t\in T$ is used to represent any token that starts a labelled span, a single tag $I$ for tokens that continue a labelled span and $O$ to tag tokens outside labelled spans.
@@ 4 Experiments
This section describes our experiments for Sequence Labelling using the HyperLoader model.
@@ 4.1 Corpora
We use seven publicly available corpora for Sequence Labelling tasks (slot-filling and Named Entity Recognition) to cover different domains and distributions and gather robust empirical evidence.
For slot-filling, we use four different dialogue-oriented datasets. First, the ATIS dataset Hemphill et al. (1990) consists of manual transcripts of audio recordings about people asking for flight information on automated airline travel inquiry systems. Second, the SNIPS dataset Coucke et al. (2018) comprises users’ intent queries distributed in seven domains: search creative work, get weather conditions, restaurant reservations, play music, add elements to a playlist, and search screening events. The third and fourth datasets are the English portions of two multilingual task-oriented dialogue datasets: The mTOP dataset Li et al. (2021) consists of eleven domains: alarm, calling, event, messaging, music, news, people, recipes, remainders, timer and weather conditions. Finally, the mTOD dataset Schuster et al. (2019) contains nine labels across three domains: alarm, reminders and weather conditions.
For Named Entity Recognition, we use the MIT Corpora (Movie, MovieTrivia and Restaurant) [^1] Downloaded from https://groups.csail.mit.edu/sls/. In the case of Movie and MovieTrivia datasets, both are composed of 12 labels, and the latter contains more complex queries. Finally, the Restaurant dataset contains 8 labels such as restaurant names, ratings, dishes and opening times. We show the complete list of labels for each dataset in Appendix A. Based on the label’s name, we obtained the label overlap across the datasets. Movie and MovieTrivia have the highest number of overlaps with four labels: genre, year, plot, director, and actor. These datasets also share the labels "genre" and "year" with the SNIPS dataset. The mTOD, mTOP and restaurant datasets share the label "location". Movie and Restaurant datasets share the label "rating". Finally, Restaurant and SNIPS datasets share two labels: "restaurant_name" and "cuisine". Nine different labels are shared across the datasets, representing 3.78% of the total of labels.
@@
Table 2: Statistics per partition of the used datasets.
4.2 Experimental Details and Evaluation
We conduct experiments to compare our model with other parameter-efficient fine-tuning methods. Our first goal is to assess the advantages of combining different parameter-efficient techniques in a multi-task setting versus a single-task setting. The second goal is to evaluate our model against other hypernetwork-based approaches in a multi-task environment.
We compare our approaches with the UniPELT framework Mao et al. (2022), HyperFormer model Karimi Mahabadi et al. (2021b) and Hyperdecoder model Ivison and Peters (2022) described in section 2. We also compare our approach using Adapters and LoRA with T5 as the backbone model in a single-task setting. We found that the prefix-tuning and the gating mechanism proposed in the UniPELT framework decrease the performance using T5. Therefore, we only use Adapters and LoRA without a gating mechanism for our baseline comparison for the combination of UniPELT and T5. For all the implementations, we use the HuggingFace’s Transformers Wolf et al. (2020) and AdapterHub Poth et al. (2023) libraries.
@@ 5.1 Full Dataset Performance
Table 3(a) shows the performance of our proposed approach and the baselines using the 100% of training and evaluation data. The UniPELT framework, Adapters and LoRA follow a single-task fine-tuning approach, therefore, the trainable parameters correspond to each dataset considered. For UniPELT, we report the results using BERT as in the original paper and T5 to compare the performance depending on the model’s architecture. Initial experiments show that the use of the Prefix-tuning method and the gating mechanism decreases the performance of the encoder-decoder model; for this reason, we do not use these elements when applying UniPELT to T5.
@@
Based on the results in Table 3(a), it is possible to observe that our proposed HyperLoader model outperforms the baselines in three of the seven datasets and obtains the highest average with 0.8811, followed by the HyperFormer model (0.8795), UniPELT with T5 (0.8740), T5 with Adapters (0.8735), T5 with LoRA (0.8717), the Hyperdecoder model (0.8709) and UniPELT with BERT (0.8537). In terms of single-task and multi-task settings, it is possible to observe that the combination of UNiPELT with T5 outperforms the other models in two datasets (Movie with 0.8839, and ATIS with 0.9611); this result demonstrates that the use of hypernetworks improves the sharing information capacity of the models while reducing the negative transfer between tasks. Finally, when comparing the multi-task approaches based on hypernetworks, HyperLoader obtained a better average result than the HyperFormer model, which is the primary foundation of our work. In addition, since Hyperdecoder has more trainable parameters, the results indicate that the superior performance of our model is not only due to having more trainable parameters than the HyperFormer approach but also due to the combination of different parameter-efficient methods in a multi-task setting.
Adapters LoRA UniPELT UniPELT HyperFormer HyperLoader Hyperdecoder
Dataset T5[ [^2] M/3.5M] $\clubsuit$ T5[ [^2] M/0.84M] $\clubsuit$ BERT[110M/1.8M] $\clubsuit$ T5[220M/4.5M] $\clubsuit$ T5[228M/5M] T5[229M/6M] T5[239M/16M]
MovieTriva 0.7184 0.7183 0.6873 0.7210 0.7295 0.7140 0.7072
Movie 0.8793 0.8709 0.8500 0.8839 0.8788 0.8745 0.8700
Restaurant 0.8069 0.8186 0.7689 0.8039 0.8136 0.7965 0.7891
ATIS 0.9581 0.9549 0.9446 0.9611 0.9595 0.9599 0.9519
SNIPS 0.9573 0.9506 0.9238 0.9478 0.9462 0.9592 0.9377
mTOP 0.8299 0.8254 0.8432 0.8374 0.8638 0.8983 0.8752
mTOD 0.9647 0.9633 0.9580 0.9627 0.9648 0.9652 0.9649
Average 0.8735 0.8717 0.8537 0.8768 0.8795 0.8811 0.8709
(a) Full dataset performance
Adapters LoRA UniPELT UniPELT HyperFormer HyperLoader Hyperdecoder
Dataset T5[ [^2] M/3.5M] $\clubsuit$ T5[ [^2] M/0.84M] $\clubsuit$ BERT[110M/1.8M] $\clubsuit$ T5[220M/4.5M] $\clubsuit$ T5[228M/5M] T5[229M/6M] T5[239M/16M]
10% of Training and validation data
MovieTriva 0.6633 0.6695 0.5152 0.6717 0.6589 0.6695 0.6597
Movie 0.8281 0.8134 0.7809 0.8244 0.8311 0.8368 0.8202
Restaurant 0.7292 0.7171 0.5907 0.7049 0.7389 0.7499 0.7216
ATIS 0.9129 0.8329 0.5901 0.8919 0.9162 0.9172 0.9193
SNIPS 0.8647 0.8460 0.7398 0.8991 0.8943 0.8970 0.8727
mTOP 0.7559 0.6886 0.6048 0.7467 0.7976 0.8114 0.7760
mTOD 0.9459 0.9453 0.9351 0.9436 0.9467 0.9500 0.9436
Average 0.8143 0.7875 0.6795 0.8117 0.8262 0.8331 0.8162
20% of Training and validation data
MovieTriva 0.6896 0.6934 0.6105 0.7036 0.6729 0.6863 0.6767
Movie 0.8492 0.8276 0.8047 0.8444 0.8407 0.8546 0.8377
Restaurant 0.7699 0.7522 0.7114 0.7562 0.7517 0.7601 0.7401
ATIS 0.9295 0.9222 0.8048 0.9281 0.9382 0.9401 0.9311
SNIPS 0.9250 0.9096 0.8474 0.9235 0.9252 0.9238 0.9044
mTOP 0.7798 0.7597 0.7052 0.7863 0.8399 0.8474 0.8140
mTOD 0.9424 0.9543 0.9433 0.9522 0.9543 0.9541 0.9529
Average 0.8408 0.8313 0.7753 0.8420 0.8461 0.8523 0.8367
(b) Low-resource setting results
@@
Table 3: Performance of the HyperLoader model and baselines is shown for 100% (a), 10%, and 20% (b) of the training and evaluation data. Each method includes the model used and the total/trainable parameters in brackets. $\clubsuit$ denotes a single-task fine-tuning approach, so trainable parameters are specific to each dataset. Bold fonts highlight the best result for each dataset and on average. Micro-averaged F1-score is reported.
@@ 5.2 Low-resource Setting Results
We evaluate the performance of our proposed model and the baselines in a low-resource setting to measure its robustness with a data scarcity constraint, a common scenario in NLP Hedderich et al. (2021). For this purpose, we down-sampled the training and validation partitions of all datasets to 10% and 20% of their original size and evaluated them in the full test subset. Table 3(b) shows the results of these experiments.
@@
The first part of Table 3(b) contains the performance of the HyperLoader model and the baselines when only 10% of the training and validation data is used. It is evident that the combination of the UniPELT framework and the BERT model experiences the most significant performance decrease, from an average of 0.8537 to 0.6795. In contrast, the T5-based approaches show a smaller performance decrease in both single-task and multi-task settings. These results suggest that an encoder-decoder model is more capable of handling low-resource settings in both single-task and multi-task configurations. For the multi-task approaches, our HyperLoader model outperforms the HyperFormer and Hyperdecoder approaches in six out of seven datasets and on the average performance with a score of 0.8331, followed by HyperFormer (0.8262) and Hyperdecoder (0.8161).
@@
The second part of Table 3(b) shows the results when 20% of the training and evaluation data is used. Using this amount of data our HyperLoader model outperforms the baselines in three of seven datasets and on the average performance. It is also noticeable that the combination of UniPELT and T5 using a single-task fine-tuning approach outperforms the other models only in the MovieTrivia dataset, with a micro-averaged F1-score of 0.7036 followed by HyperLoader with 0.6863, Hyperdecoder with 0.6767, HyperFormer with 0.6729 and finally UniPELT using BERT with 0.6105. These results suggest that combining various parameter-efficient methods in a multi-task setting and generating their weights based on the task, transformer layer, and the position where the parameter-efficient method is applied (Adapter layer or LoRA matrix) is effective for Sequence Labelling tasks, leading to good performance with a parameter-efficient approach. Figure 2 shows the relationship between average performance and the percentage of trainable parameters per model. The symbol $\clubsuit$ denotes a single fine-tuning approach. Our approach consistently outperforms other methods across 10%, 20%, and 100% of the training and validation data. This evidence shows that combining various parameter-efficient fine-tuning techniques in a multi-task setting is effective for low-resource scenarios. The performance boost is not just due to adding more parameters, as seen with the hyperdecoder model and the combination of UniPELT with BERT, which have a higher percentage of trainable parameters. Therefore, the combination of parameter-efficient methods and weight generation is a potent alternative for Sequence Labelling tasks.
@@
Figure 2: Average performance-percentage of trainable parameters plot using different portions of the datasets. $\clubsuit$ indicates a single-task fine-tuning approach.
5.3 Ablation Experiments
We conducted ablation experiments to assess the effect of varying the number of trainable parameters in our model. Table 4 presents the results in descending order of performance. The first row (in blue) corresponds to the final version of our model, as described in section 3. The fifth row (in grey) corresponds to the HyperFormer model, which is the main foundation of our work.
The initial experiments (denoted by +) involved increasing the number of trainable parameters for generating LoRA’s weight matrices. In the case of using four hypernetworks, the first one generates the Low-rank A matrix for Query (Q) and Value (V) in the self-attention module, and the second generates the Low-rank B matrix for Q and V in the self-attention module. Analogously, the third and fourth hypernetworks generate the A and B matrices in the cross-attention module of the decoder. In the case of 8 hypernetworks, each generates the weight of only one component: one for the A and one for the B matrices in the Query; another two hypernetworks for the A and B matrices in the Value of the self-attention. The remaining four hypernetworks are analogous to the former but in the cross-attention component of the decoder. Finally, the last experiment involves increasing the reduction factor ($r=16$) of the LoRA matrices in the final version of the HyperLoader model. These results indicate that our model achieves a strong balance between performance and the percentage of trainable parameters (2.9543%), with only a 0.6694% increase compared to the HyperFormer model (2.3849%). It also shows that using only 2 hypernetworks for the LoRA matrices enables the model to generate better weights based on the task, the parameter-efficient method and its position within the Transformer layer of the model.
The second experiment involved removing (denoted by --) the Adapters from the model and adjusting only the task-conditioned LoRA matrices during training. This version showed the lowest average performance, though the difference from the HyperFormer model was minimal (0.0001). This suggests that while a single parameter-efficient technique can effectively address different tasks, combining such methods provides better results in a multi-task setting and shows that further research into this area is necessary.
Variant Total/trainable params Avg score
HyperLoader 229M/6M 0.8811
+ 4 LoRA HNs 228M/8M 0.0808
+ 8 LoRA HNs 234M/11M 0.800
+ 2 LoRA HNs, r=16 229M/6M 0.8804
HyperFormer 228M/5M 0.8795
- Adapters 225M/1.8M 0.8794
Table 4: Average performance of the ablation experiments. We denote hypernetworks as HNs.